# Learning Flexible Body Collision Dynamics with Hierarchical Contact Mesh Transformer

### v1

@@ Abstract
Recently, many mesh-based graph neural network (GNN) models have been proposed for modeling complex high-dimensional physical systems. Remarkable achievements have been made in significantly reducing the solving time compared to traditional numerical solvers. These methods are typically designed to i) reduce the computational cost in solving physical dynamics and/or ii) propose techniques to enhance the solution accuracy in fluid and rigid body dynamics. However, it remains under-explored whether they are effective in addressing the challenges of flexible body dynamics, where instantaneous collisions occur within a very short timeframe. In this paper, we present Hierarchical Contact Mesh Transformer (HCMT), which uses hierarchical mesh structures and can learn long-range dependencies (occurred by collisions) among spatially distant positions of a body — two close positions in a higher-level mesh corresponds to two distant positions in a lower-level mesh. HCMT enables long-range interactions, and the hierarchical mesh structure quickly propagates collision effects to faraway positions. To this end, it consists of a contact mesh Transformer and a hierarchical mesh Transformer (CMT and HMT, respectively). Lastly, we propose a flexible body dynamics dataset, consisting of trajectories that reflect experimental settings frequently used in the display industry for product designs. We also compare the performance of several baselines using well-known benchmark datasets. Our results show that HCMT provides significant performance improvements over existing methods.
@@ 1 Introduction
Figure 1: The relationship between velocity and non-linearity in various physical systems. Typically, implicit methods are used in the static system, while explicit methods (Chang, 2007) are employed in the dynamics domain. Flexible dynamics solve problems that have highly non-linear characteristics and occur quickly in a very short time.
Figure 1: The relationship between velocity and non-linearity in various physical systems. Typically, implicit methods are used in the static system, while explicit methods (Chang, 2007) are employed in the dynamics domain. Flexible dynamics solve problems that have highly non-linear characteristics and occur quickly in a very short time.
Figure 2: An illustration of the contact propagation via one-hop propagation in various levels. Level $i$ is the $i$-th level mesh (after pooling nodes from the previous level) in our hierarchical mesh structure. After the node pooling, our Transformer accelerates the contact propagation, which is a key in simulating flexible body collision dynamics.
@@
For addressing these challenges, we design Hierarchical Contact Mesh Graph Transformer (HCMT). The multi-level hierarchical structure of HCMT i) quickly propagates collisions and ii) decreases training time due to the reduced number of nodes in Level $i$, i.e., the $i$-th mesh, resulting from pooling nodes from previous Level (cf. Fig. 2). HCMT with higher-level meshes focuses on long-range dynamic interactions. In addition, HCMT has two Transformers: one for contact dynamics, and the other for flexible dynamics. Lastly, we introduce a novel benchmark dataset named Impact Plate. Impact Plate replicates simulations conducted in the display industry for mobile phone display rigidity assessments. We evaluate our model on three publicly available datasets (Sphere Simple (Pfaff et al., 2020), Deforming Plate (Pfaff et al., 2020), Deformable Plate (Linkerhägner et al., 2023)) and the novel Impact Plate, and our model achieves consistently the best performance.
@@
  • We propose Hierarchical Contact Mesh Transformer (HCMT), which, to the best of our knowledge, incorporates collisions into flexible body dynamics for the first time.
@@
  • HCMT outperforms other baseline models in static, rigid dynamics, and flexible dynamics systems.
@@ 2.1 GNN-Based Models and Hierarchical Models in physical systems
Recently, several hierarchical models have been introduced to increase the propagation radius (Gao & Ji, 2019; Fortunato et al., 2022; Han et al., 2022; Janny et al., 2023; Cao et al., 2022; Grigorev et al., 2023). Hierarchical models in these studies are divided into two types: The first type is a dual-level structure: i) GMR-Transformer-GMUS (Han et al., 2022) uses a pooling method to select pivotal nodes through uniform sampling, ii) EAGLE Transformer (Janny et al., 2023) proposes a clustering-based pooling method and shows promising performance in fluid dynamics, and iii) MS-MGN (Fortunato et al., 2022) proposes a dual-layer framework that performs message passing at two different resolutions (fine and coarse resolutions) for mesh-based simulation learning. The second type has a multi-level structure: i) Cao et al. (2022) analyzes the limitations of existing pooling strategies and proposes bi-stride pooling, which uses breadth-first search (BFS) to select nodes, ii) HOOD (Grigorev et al., 2023) leverages multi-level message passing with unsupervised training to predict clothing dynamics in real-time for arbitrary types of garments and body shapes.
@@ 3.3 Learning Flexible Body Dynamics
We discuss based on Impact Plate. According to the MGN approach, the model predicts one or more output features $\hat{\mathbf{o}}_{i}$ such as velocity $\hat{\dot{\mathbf{x}}}^{t}_{i}$ and next stress $\hat{\mathbf{\sigma}}^{t+1}_{i}$ by employing an MLP decoder. Finally, $\hat{\dot{\mathbf{x}}}^{t}_{i}$ is used to calculate the next position $\hat{\mathbf{x}}^{t+1}_{i}$ through updater, which performs a first-order integration ($\hat{\mathbf{x}}^{t+1}_{i}=\hat{\dot{\mathbf{x}}}^{t}_{i}+\mathbf{x}^{t}_{i}$).
@@
We use the one-step MSE loss as the training objective. Since stresses are included, the position and stress MSE in flexible dynamics are calculated as follows:
@@ 3.4 Discussion
EAGLE (Janny et al., 2023), which learns fluid dynamics, does not use edge features in its Transformer but in its last GNN. In Eulerian-based fluid dynamics, node positions and edge lengths are fixed, making edge functions less useful. In flexible dynamics where contact occurs, however, contact edges carry important messages. In addition, because the object is severely deformed, mesh edge features are also greatly affected. GMR-Transformer-GMUS (Han et al., 2022) is also a representative example of applying a Transformer in the mesh domain. GMR-Transformer-GMUS uses one Transformer without edge features, while our model uses two Transformers with edge features. Graph Transformer (Dwivedi & Bresson, 2020) (GT) has been extended to incorporate edge representations, which can work with information associated with edges. In comparison with them, we utilize both mesh and contact edge features to infer accurately.
@@ 4.1 Experimental Settings
As baselines, we use MGN (Pfaff et al., 2020), a state-of-the-art model in the field of complex physics, and Graph Transformer (GT) (Dwivedi & Bresson, 2020) with edge features. Detailed descriptions for all the baselines, hardware and software specifications are provided in Appendix D. All experiments performed on NVIDIA 3090 and Intel Core-i9 CPUs. The number of blocks $L=L_{C}+L_{H}$ is set to 15. For further details on hyperparameters, see Table 8 in Appendix E.
@@ 4.2 Experimental Results
Comparison to MGN and GT
HCMT shows superior performance in terms of long-term predictions across all domains compared to MGN and GT. In particular, our method significantly outperforms others in Impact Plate, which incurs strong impacts. Our hierarchical architecture propagates collisions in space more efficiently and effectively than others. Fig. 6 shows that as the level increases, shortcut message passing is possible in our method. Fig. 7 is a contour image showing the stress on the falling ball and plate after a collision, visually demonstrating its superiority. HCMT can be generalized for unseen test cases. We report the generalization abilities of HCMT in Appendix G.
@@
Figure 6: Visualization of meshes by the level on Impact Plate. As the level increases, the distance between neighboring nodes increases, allowing for long-range interactions within each body via HMT. See more visualizations in Appendix B.
@@
Figure 7: 2D cross-sectional contour of the stress field in Impact Plate. In the red bounding box, HCMT is the most similar to the node positions in the ground truth. A brighter color means a higher stress. Rollout images of other datasets can be found in Appendix L.
@@
We visualize the dual-branch self-attention maps of CMT layer on Impact Plate in Fig. 8 — we note that the self-attention is naturally sparse in $G_{0}$. The contact and mesh self-attention show non-overlapping attention maps and shows the varying importance of contact and mesh edges. The red bounding box in Fig. 8 (a) represents the importance of contact edges between the ball and the plate, and the blue box in Fig. 8 (b) signifies the self-attention among plate nodes, and the yellow box represents interactions through edges among ball nodes (See Appendix J for self-attention map of HMT and results in different datasets).
@@ 4.3 Ablation and Sensitivity Studies
Table 3: The results of ablation studies. The values corresponding to each dataset represent the RMSE (rollout-all, ×$\times 10^{3}$). bold denotes the best performance.
@@ Appendix B Pooling Method Details
Max Angle (${}^{\circ}$) Min Angle (${}^{\circ}$) Jacobian Max Angle (${}^{\circ}$) Min Angle (${}^{\circ}$) Jacobian

### v3

@@ Abstract
Recently, many mesh-based graph neural network (GNN) models have been proposed for modeling complex high-dimensional physical systems. Remarkable achievements have been made in significantly reducing the solving time compared to traditional numerical solvers. These methods are typically designed to i) reduce the computational cost in solving physical dynamics and/or ii) propose techniques to enhance the solution accuracy in fluid and rigid body dynamics. However, it remains under-explored whether they are effective in addressing the challenges of flexible body dynamics, where instantaneous collisions occur within a very short timeframe. In this paper, we present Hierarchical Contact Mesh Transformer (HCMT), which uses hierarchical mesh structures and can learn long-range dependencies (occurred by collisions) among spatially distant positions of a body — two close positions in a higher-level mesh correspond to two distant positions in a lower-level mesh. HCMT enables long-range interactions, and the hierarchical mesh structure quickly propagates collision effects to faraway positions. To this end, it consists of a contact mesh Transformer and a hierarchical mesh Transformer (CMT and HMT, respectively). Lastly, we propose a flexible body dynamics dataset, consisting of trajectories that reflect experimental settings frequently used in the display industry for product designs. We also compare the performance of several baselines using well-known benchmark datasets. Our results show that HCMT provides significant performance improvements over existing methods. Our code is available at https://github.com/yuyudeep/hcmt.
@@ 1 Introduction
Figure 1: Relationship between velocity and non-linearity in various physical systems. Typically, implicit methods are used in static system, while explicit methods (Chang, 2007) are employed in dynamic domain. Flexible dynamics solve problems with highly non-linear characteristics and occur quickly in a very short time.
Figure 2: An illustration of the contact propagation via one-hop propagation in various levels. Level $i$ is the $i$-th level mesh (after pooling nodes from the previous level) in our hierarchical mesh structure. After node pooling, our Transformer accelerates contact propagation, which is a key in simulating flexible body collision dynamics.
@@
To address these challenges, we propose a novel method called Hierarchical Contact Mesh Graph Transformer (HCMT). The multi-level hierarchical structure of HCMT i) quickly propagates collisions and ii) decreases training time due to the reduced number of nodes in Level $i$, i.e., the $i$-th mesh, resulting from pooling nodes from previous level (cf. Fig. 2). HCMT with higher-level meshes focuses on long-range dynamic interactions. In addition, HCMT has two Transformers: one for contact dynamics, and the other for flexible dynamics. Finally, we introduce a novel benchmark dataset named Impact Plate. Impact Plate replicates simulations conducted in the display industry for mobile phone display rigidity assessments. We evaluate our model on three publicly available datasets (Sphere Simple (Pfaff et al., 2020), Deforming Plate (Pfaff et al., 2020), Deformable Plate (Linkerhägner et al., 2023)) and the novel Impact Plate, and our model achieves consistently the best performance.
@@
  • We propose a Hierarchical Contact Mesh Transformer (HCMT), which, to the best of our knowledge, incorporates collisions into flexible body dynamics for the first time.
@@
  • HCMT outperforms baseline models in static, rigid, and flexible dynamics systems.
@@ 2.1 GNN-Based Models and Hierarchical Models in physical systems
Recently, several hierarchical models have been introduced to increase the propagation radius (Gao & Ji, 2019; Fortunato et al., 2022; Han et al., 2022; Janny et al., 2023; Cao et al., 2022; Grigorev et al., 2023). Hierarchical models in these studies are divided into two types: The first type is a dual-level structure: i) GMR-Transformer-GMUS (Han et al., 2022) uses a pooling method to select pivotal nodes through uniform sampling, ii) EAGLE Transformer (Janny et al., 2023) proposes a clustering-based pooling method and shows promising performance in fluid dynamics, and iii) MS-MGN (Fortunato et al., 2022) proposes a dual-layer framework that passes messages on two different resolutions (fine and coarse resolutions) for mesh-based simulation learning. The second type has a multi-level structure: i) Cao et al. (2022) analyzes the limitations of existing pooling strategies and proposes bi-stride pooling, which uses breadth-first search (BFS) to select nodes, ii) HOOD (Grigorev et al., 2023) leverages multi-level message passing with unsupervised training to predict clothing dynamics in real-time for arbitrary types of garments and body shapes.
@@ 3.3 Learning Flexible Body Dynamics
We describe the decoder and updater based on Impact Plate dataset. According to the MGN approach, the model predicts one or more output features $\hat{\mathbf{o}}_{i}$ such as velocity $\hat{\dot{\mathbf{x}}}^{t}_{i}$ and next stress $\hat{\mathbf{\sigma}}^{t+1}_{i}$ by employing an MLP decoder. Finally, $\hat{\dot{\mathbf{x}}}^{t}_{i}$ is used to calculate the next position $\hat{\mathbf{x}}^{t+1}_{i}$ through the updater, which performs a first-order integration ($\hat{\mathbf{x}}^{t+1}_{i}=\hat{\dot{\mathbf{x}}}^{t}_{i}+\mathbf{x}^{t}_{i}$).
@@
We use the one-step MSE loss as a training objective. Since stresses are included, the position and stress MSE in flexible dynamics are calculated as follows:
@@ 3.4 Discussion
EAGLE (Janny et al., 2023), which learns fluid dynamics, does not use edge features in its Transformer but in its last GNN. In Eulerian-based fluid dynamics, node positions and edge lengths are fixed, making edge functions less useful. In flexible dynamics where contact occurs, however, contact edges carry important messages. In addition, because the object is severely deformed, mesh edge features are also greatly affected. GMR-Transformer-GMUS (Han et al., 2022) is a representative example of applying a Transformer in the mesh domain. GMR-Transformer-GMUS uses one Transformer without edge features, while our model uses two Transformers with edge features. Graph Transformer (Dwivedi & Bresson, 2020) (GT) has been extended to incorporate edge representations, which can work with information associated with edges. In comparison with them, we utilize both mesh and contact edge features to infer accurately.
@@ 4.1 Experimental Settings
As baselines, we use MGN (Pfaff et al., 2020), a state-of-the-art model in the field of complex physics, and GT (Dwivedi & Bresson, 2020) with edge features. A detailed description for all baselines is provided in Appendix D. The number of blocks $L=L_{C}+L_{H}$ is set to 15. For further details on hyperparameters, see Table 8 in Appendix E.
@@ 4.2 Experimental Results
Figure 6: Level-wise mesh visualization of Impact Plate. See more visualizations in Appendix B.
@@
Figure 7: 2D cross-sectional contour of the stress field in Impact Plate. In the red bounding box, HCMT is the most similar to the node positions in the ground truth. Brighter colors mean higher stress. Rollout images of other datasets can be found in Appendix L.
Result
HCMT shows superior performance in terms of long-term prediction across all domains compared to MGN and GT. In particular, we can see that HCMT outperforms other methods in Impact Plate, where strong impacts occur, and that our hierarchical architecture effectively propagates collisions. Fig. 6 shows that as the level increases, shortcut message passing is possible in our method. Fig. 7 shows the stress on the falling ball and plate after a collision, visually demonstrating the superiority of HCMT. We also report generalization abilities of HCMT in Appendix G.
@@
We visualize the dual-branch self-attention maps of CMT layer on Impact Plate in Fig. 8 — we note that the self-attention is naturally sparse in $G_{0}$. The contact and mesh self-attention show non-overlapping attention maps and the varying importance of contact and mesh edges. The red bounding box in Fig. 8 (a) represents the importance of contact edges between the ball and the plate, and the blue box in Fig. 8 (b) signifies the self-attention among plate nodes, and the yellow box represents interactions through edges among ball nodes (see Appendix J for self-attention map of HMT and results in different datasets).
@@ 4.3 Ablation and Sensitivity Studies
Table 3: The results of ablation studies.
@@ Appendix B Pooling Method Details
Max Angle () Min Angle () Jacobian Max Angle () Min Angle () Jacobian